OpenStudy (anonymous):

3. 0.5b + 4 = 2(b + 2) (1 point

OpenStudy (anonymous):

0.5b+4=2b+4 0.5b=2b b=4

OpenStudy (anonymous):

4. 8 – (3 + b) = b – 9 (1 point